Solve the Proportion:
5/p+1 = 4p/p

Answer by checkley77(12844)

5/p+1 = 4p/p
5/p+1=4
5/p=4-1
5/p=3
3p=5
p=5/3
ans.
Proof:
5/(5/3)+1=4(5/3)/(5/3)
5*3/5+1=4(5/3*3/5)
3+1=4
4=4